OpenStudy (anonymous):

10ac-x -------=-3, for a 11

OpenStudy (anonymous):

10ac-x=-33 ac=(-33+x)/10 a=((-33+x)/10)/c a=c*(-33+x)/10

OpenStudy (anonymous):

a=(-33c+xc)/10